# On the Expressiveness of Visual Prompt Experts

### v5

@@ 7 Conclusion
Supplement to “On the Expressiveness of Visual Prompt Experts”
In this supplementary material, we provide detailed proofs of the main results in Appendix A and additional theoretical findings in Appendix B. A discussion of related work is presented in Appendix C. Implementation details of our experiments are described in Appendix D, while additional experimental results are included in Appendix E. Specifically, these include detailed per-task results corresponding to the main experiments in Section 6 (Appendices E.1, E.2), statistical significance tests (Appendix E.3), an evaluation of VAPT performance across different backbone scales (Appendix E.4), and semantic segmentation results (Appendix E.5). Finally, ablation studies, computational cost analyses, and interpretive visualizations are presented in Appendix E.6, Appendix E.7, and Appendix E.8, respectively.
@@ A.1.1 Local Part
$+\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}})\big{[}\sigma^{(1)}(W_{*,1j}{\bm{X}})\sigma(W_{*,1j}{\bm{X}}){\bm{X}}^{\top}L_{n,1,j}{\bm{X}}W_{*,2}$
@@
$-\sum_{j:|\mathcal{V}_{j}|=1}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}})\big{[}\sigma^{(1)}(W_{*,1j}{\bm{X}}){\bm{X}}^{\top}L_{n,1,j}{\bm{X}}+\sigma(W_{*,1j}{\bm{X}})L_{n,2,j}^{\top}{\bm{X}}\big{]}f_{G_{n}}({\bm{X}})$
$-\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}})\big{[}\sigma^{(1)}(W_{*,1j}{\bm{X}}){\bm{X}}^{\top}L_{n,1,j}{\bm{X}}+\sigma(W_{*,1j}{\bm{X}})L_{n,2,j}^{\top}{\bm{X}}$
@@
$-\sum_{j=1}^{L}M_{n,j,\mathbf{0}_{d},\mathbf{0}_{d}}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}})f_{G_{n}}({\bm{X}})$
$+\sum_{j=1}^{L}M_{n,j,\mathbf{0}_{d},\mathbf{0}_{d}}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}}){\sigma}(W_{*,1j}{\bm{X}})W_{*,2}$
$+\tilde{R}_{n,1}({\bm{X}})+\tilde{R}_{n,2}({\bm{X}})-\bar{R}_{n,3}({\bm{X}})-\bar{R}_{n,4}({\bm{X}})$ (22)
@@ A.1.2 Global Part
$\Bigg{\{}\frac{\exp((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}})}{D_{f,G}({\bm{X}})}:j^{\prime}\in[L]\Bigg{\}}$ =$=\Bigg{\{}\frac{\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})}{D_{f,G_{*}}({\bm{X}})}:j^{\prime}\in[L]\Bigg{\}},$
@@
$=\sum_{j^{\prime}=1}^{L}\exp{(b_{*,j})}\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}})CW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}),$ (28)
@@ A.2 Proof of Proposition 1
$+\sum_{j^{\prime}=1}^{L}\frac{\exp((BW^{(2)}_{*}\sigma(W^{(1)}_{*,j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})}{D_{f}({\bm{X}})}\cdot CW^{(2)}_{*}\sigma(W^{(1)}_{*,j^{\prime}}{\bm{X}})$
@@
where $p(Y_{i}|f_{G}({\bm{X}}_{i}),\sigma^{2}I_{d^{\prime}})$ is the multivariate Gaussian distribution with mean $f_{G}({\bm{X}})$ and covariance matrix $\sigma^{2}I_{d^{\prime}}$. From the result of Theorem 7.4 in [68], it follows that
@@ Appendix B Additional Theoretical Results for Visual Adaptive Prompt Tuning
where $\varepsilon_{1},\ldots,\varepsilon_{n}$ are independent Gaussian noise variables such that $\mathbb{E}[{\varepsilon_{i}}|{\bm{X}}_{i}]=0$ and $\mathrm{Var}(\varepsilon_{i}|{\bm{X}}_{i})=\nu^{2}I_{d^{\prime}}$ for all $1\leq i\leq n$. Additionally, we assume that ${\bm{X}}_{1},{\bm{X}}_{2},\ldots,{\bm{X}}_{n}$ are i.i.d. samples from some probability distribution $\mu$. The regression function $f_{\widetilde{G}_{*}}(\cdot)$ in Equation (29) then takes the form
@@ B.0.1 Local Part
$0=\lim_{n\to\infty}\dfrac{\|f_{G_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}}{m_{n}\mathcal{D}_{2n}}\geq\int\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{\widetilde{G}_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}d\mu({\bm{X}})\geq 0$
@@ B.0.2 Global Part
$\Big{\{}\frac{\exp((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}})}{\tilde{D}_{f,G}({\bm{X}})}:j^{\prime}\in[L]\Big{\}}$ =$=\Big{\{}\frac{\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})}{\tilde{D}_{f,\widetilde{G}_{*}}({\bm{X}})}:j^{\prime}\in[L]\Big{\}},$
@@
$=\sum_{j^{\prime}=1}^{L}\exp{(b_{*,j})}\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}})CW_{*,2}W_{*,1j^{\prime}}{\bm{X}},$ (40)
@@ Appendix C Related Work
Mixture of Experts. Building on classical mixture models with adaptive gating mechanisms [30, 33, 77], the MoE model has been extensively studied and refined over the years. Notably, subsequent works [12, 65] introduced the MoE layer as an efficient tool for scaling model capacity. Unlike traditional models that apply uniform parameters to all inputs, the MoE layer applies specific parameter subsets for each input, creating a sparsely activated layer that enables significant capacity growth without proportional increase in computational cost [13, 86]. This efficiency and scalability have driven its adoption across diverse domains and tasks [63, 10, 66].
@@
Theory of Mixture of Experts. Although MoEs have been widely used to scale up large models, their theoretical foundations remain under active development. For example, [24] focused on input-free gating Gaussian MoEs and showed that under maximum likelihood estimation, the experts’ convergence rates depend on the algebraic independence of the expert functions. Next, [53, 50] established convergence rates for both density and parameter estimation in Softmax-gating Gaussian MoEs, linking these rates to the solvability of polynomial systems under Voronoi-based loss functions. More recently, [51, 52] employed least-squares estimation to identify conditions under which expert functions are identifiable. Under these conditions, the resulting estimation rates improve substantially.
@@
Parameter-Efficient Fine-Tuning. Fine-tuning pre-trained foundational models [9, 45, 36] has become a widely adopted strategy for tackling downstream tasks [76]. While effective, full fine-tuning is resource-intensive, requiring updates to all network parameters and the storage of a separate fine-tuned model for each task [19]. To address these challenges, researchers have increasingly focused on parameter-efficient fine-tuning (PEFT) methods. These methods can be categorized into partial tuning, extra module, and prompt tuning. Partial tuning freezes most of the backbone, fine-tuning only a subset, such as linear heads or a few layers [46, 4, 21]. Extra module methods add trainable parameters to the backbone, such as side structures [83], residual MLP modules in Transformer layers [25, 2], or low-rank weight updates [26]. Despite their promise, partial tuning and extra module methods often face limitations that restrict their applicability. First, they may fail to achieve performance comparable to full fine-tuning [46, 4, 31, 22]. Second, some methods rely on specific architectural modifications [62, 2, 83], limiting their generalizability across different backbone architectures.
@@
Prompting Methods. Prompt tuning [44], initially proposed for language tasks, offers a simpler yet effective alternative by introducing learnable parameters to the input sequence of backbone models, updating only these parameters during fine-tuning [44, 31, 40, 38]. Despite its simplicity, prompt tuning has demonstrated significant performance gains. Recently, visual prompt tuning has emerged as a promising paradigm in PEFT techniques for computer vision. Current advancements in visual prompt tuning focus on engineering improvements, such as minimizing parameter usage [18] and broadening applicability to diverse tasks [78, 67, 79]. However, the theoretical foundations of prompt-based methods remain underexplored. For instance, [20] investigated the relationship between prompt tuning and adapter methods, while [39] analyzed these techniques within the context of MoE models. Furthermore, [60] highlighted the limitations of prompting, showing that it cannot alter relative attention patterns and instead biases attention layer outputs in a fixed direction.
Adaptive Prompting. Prior studies have explored different adaptive prompting strategies. For example, [28, 35] integrated VPT with meta-learning by clustering the downstream dataset and assigning cluster-specific prompts based on the initial ViT layer’s input ${\bm{X}}^{(0)}$. Our approach, VAPT, distinguishes itself by generating prompts at each layer based on the layer-specific input ${\bm{X}}^{(l)}$. This layer-wise conditioning aligns with our theoretical framework linking MoE and VPT, which dictates dependency on ${\bm{X}}^{(l)}$ for both experts and gating functions, facilitating rigorous analysis. Crucially, VAPT’s motivation stems from the functional properties of prompt experts and their disparity with pre-trained experts, not merely from the aim of using different prompts per input. In contrast, [54, 59, 27], add adaptive components directly to layer inputs. From our MoE perspective, this risks altering the pre-trained experts $f_{1},\dots,f_{N}$ and disrupting their learned knowledge from the pre-training process. Furthermore, this approach fixes the number of adaptive components to that of the pre-trained experts and lacks explicit theoretical grounding. VAPT, conversely, introduces new prompt experts alongside frozen pre-trained ones similar to VPT, preserving knowledge, offering flexibility in the number of experts, and is grounded on a solid theoretical foundation derived from the MoE viewpoint. Finally, while some adaptive prompting methods employ complex image-to-prompt transformations that can obscure analysis of their effectiveness, VAPT utilizes a simple yet effective formulation, enabling more straightforward and rigorous theoretical examination.
@@ Appendix D Implementation Details
Table 4: Specifications of the datasets used in our experiments, which include 24 datasets from two benchmarks: FGVC and VTAB-1K [82].
@@
CUB-200-2011 [72] Fine-grained recognition of bird species 200 5,394 600 5,794
NABirds [69] Fine-grained recognition of bird species 55 21,536 2,393 24,633
Oxford Flowers [55] Fine-grained recognition of flower species 102 1,020 1,020 6,149
Stanford Dogs [34] Fine-grained recognition of dog species 120 10,800 1,200 8,580
Stanford Cars [15] Fine-grained recognition of car 196 7,329 815 8,041
@@
CIFAR-100 [37] Natural 100 800/1000 200 10,000
Caltech101 [14] 102 6,084
DTD [6] 47 1,880
Flowers102 [55] 102 6,149
Pets [57] 37 3,669
SVHN [49] 10 26,032
Sun397 [75] 397 21,750
Patch Camelyon [71] Specialized 2 800/1000 200 32,768
EuroSAT [23] 10 5,400
Resisc45 [5] 45 6,300
Retinopathy [17] 5 42,670
Clevr/count [32] Structured 8 800/1000 200 15,000
Clevr/distance [32] 6 15,000
DMLab [1] 6 22,735
KITTI/distance [16] 4 711
dSprites/loc [48] 16 73,728
dSprites/ori [48] 16 73,728
SmallNORB/azi [42] 18 12,150
SmallNORB/ele [42] 9 12,150
Dataset Specifications. Table 4 provides a summary of the statistics and details of the classification datasets evaluated in this paper. Figure 3 presents image examples from all 24 datasets. Following [31], each FGVC dataset is randomly split into 90% train and 10% val, with the val set used for hyperparameter tuning. For VTAB-1K, we use an 800-200 split for tuning, and we train on all available data for the final evaluation.
Pre-trained Backbone Specifications. This study explores two primary groups of Vision Transformer (ViT) models. The first group comprises conventional ViT architectures [9], including ViT-Base, ViT-Large, and ViT-Huge, pre-trained on the large-scale ImageNet-21K [7] dataset. The second group focuses on self-supervised models, specifically Masked Autoencoders (MAE) [21] and Momentum Contrast v3 (MoCo v3) [4], both pre-trained on ImageNet-1K. Unless otherwise specified, we use ViT-B/16 with supervised pre-training on ImageNet-21K dataset by default.
@@
Hyperparameters. We use the val set of each dataset to find the optimal prompt length $N_{p}$, kernel size $K$, and hidden dimension $r$ of the feature projector. Following [31], the search range for $N_{p}$ is $\{1,5,10,50,100,200\}$. Given the relatively small resolution of our feature map ${\bm{X}}_{\mathrm{img}}$ (e.g., $14\times 14$ for ViT), the kernel size $K$ is selected from $\{2,3,4\}$. For the hidden dimension $r$, we explore the range $\{4,8,16,32,64,128,256\}$. To identify the optimal learning rate and weight decay, we perform a grid search, consistent with [46, 31]. The learning rate is searched within $\{50,25,10,5,2.5,1,0.5,0.25,0.1,0.05\}$, while the weight decay is chosen from $\{0.01,0.001,0.0001,0.0\}$. We adopt the batch size settings of [31], using values of 64 and 128. The models are optimized with SGD for 100 epochs, employing a cosine decay learning rate schedule with 10 warm-up epochs.
Reproducibility. Our method, VAPT, is implemented in PyTorch [58]. All experimental workflows, including training and evaluation, were conducted on NVIDIA A100-40GB GPUs. The entire implementation will be made openly available to ensure reproducibility and facilitate future research.

### v6

@@ 7 Conclusion
Reproducibility Statement
In order to facilitate the reproduction of our empirical results, we provide detailed descriptions of the experimental setup in Section 6.1 and Appendix D. All datasets used in this study are publicly available, enabling full replication of our experiments.
Supplement to “Revisit Visual Prompt Tuning: The Expressiveness of Prompt Experts”
In this supplementary material, we provide detailed proofs of the main results in Appendix A and additional theoretical findings in Appendix B. A discussion of related work is presented in Appendix C. Implementation details of our experiments are described in Appendix D, while additional experimental results are included in Appendix E. Specifically, these include detailed per-task results corresponding to the main experiments in Section 6 (Appendices E.1, E.2), statistical significance tests (Appendix E.3), an evaluation of VAPT performance across different backbone scales (Appendix E.4), and semantic segmentation results (Appendix E.5). Finally, ablation studies, computational cost analyses, adversarial robustness, and interpretive visualizations are presented in Appendix E.6, Appendix E.7, Appendix E.8, and Appendix E.9, respectively.
@@ A.1.1 Local Part
$\hskip-40.00006pt+\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}})\big[\sigma^{(1)}(W_{*,1j}{\bm{X}})\sigma(W_{*,1j}{\bm{X}}){\bm{X}}^{\top}L_{n,1,j}{\bm{X}}W_{*,2}$
@@
$\hskip-40.00006pt-\sum_{j:|\mathcal{V}_{j}|=1}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}})\big[\sigma^{(1)}(W_{*,1j}{\bm{X}}){\bm{X}}^{\top}L_{n,1,j}{\bm{X}}+\sigma(W_{*,1j}{\bm{X}})L_{n,2,j}^{\top}{\bm{X}}\big]f_{G_{n}}({\bm{X}})$
$\hskip-40.00006pt-\sum_{j:|\mathcal{V}_{j}|>1}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}})\big[\sigma^{(1)}(W_{*,1j}{\bm{X}}){\bm{X}}^{\top}L_{n,1,j}{\bm{X}}+\sigma(W_{*,1j}{\bm{X}})L_{n,2,j}^{\top}{\bm{X}}$
@@
$\hskip-40.00006pt-\sum_{j=1}^{L}M_{n,j,\mathbf{0}_{d},\mathbf{0}_{d}}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}})f_{G_{n}}({\bm{X}})$
$\hskip-40.00006pt+\sum_{j=1}^{L}M_{n,j,\mathbf{0}_{d},\mathbf{0}_{d}}\exp((W_{*,2}\sigma(W_{*,1j}{\bm{X}}))^{\top}{\bm{X}}){\sigma}(W_{*,1j}{\bm{X}})W_{*,2}$
$\hskip-40.00006pt+\tilde{R}_{n,1}({\bm{X}})+\tilde{R}_{n,2}({\bm{X}})-\bar{R}_{n,3}({\bm{X}})-\bar{R}_{n,4}({\bm{X}})$ (22)
@@ A.1.2 Global Part
$\Bigg\{\frac{\exp((BW_{2}\sigma(W_{1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{j^{\prime}})}{D_{f,G}({\bm{X}})}:j^{\prime}\in[L]\Bigg\}$ $=\Bigg\{\frac{\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})}{D_{f,G_{*}}({\bm{X}})}:j^{\prime}\in[L]\Bigg\},$
@@
$\hskip-200.0003pt=\sum_{j^{\prime}=1}^{L}\exp{(b_{*,j})}\exp((BW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}))^{\top}{\bm{X}})CW_{*,2}\sigma(W_{*,1j^{\prime}}{\bm{X}}),$ (28)
@@ A.2 Proof of Proposition 1
$\hskip-40.00006pt+\sum_{j^{\prime}=1}^{L}\frac{\exp((BW^{(2)}_{*}\sigma(W^{(1)}_{*,j^{\prime}}{\bm{X}}))^{\top}{\bm{X}}+b_{*,j^{\prime}})}{D_{f}({\bm{X}})}\cdot CW^{(2)}_{*}\sigma(W^{(1)}_{*,j^{\prime}}{\bm{X}})$
@@
where $p(Y_{i}|f_{G}({\bm{X}}_{i}),\sigma^{2}I_{d^{\prime}})$ is the multivariate Gaussian distribution with mean $f_{G}({\bm{X}})$ and covariance matrix $\sigma^{2}I_{d^{\prime}}$. From the result of Theorem 7.4 in van de Geer (2000), it follows that
@@ Appendix B Additional Theoretical Results for Visual Adaptive Prompt Tuning
[…]
@@ B.0.1 Local Part
$0=\lim_{n\to\infty}\dfrac{\|f_{G_{n}}-f_{\widetilde{G}_{*}}\|_{L_{2}(\mu)}}{m_{n}\mathcal{D}_{2n}}\geq\int\liminf_{n\to\infty}\dfrac{\left|f_{G_{n}}({\bm{X}})-f_{\widetilde{G}_{*}}({\bm{X}})\right|}{m_{n}\mathcal{D}_{2n}}d\mu({\bm{X}})\geq 0.$
@@ B.0.2 Global Part
$\Big\{\frac{\exp((BW_{2}W_{1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{j^{\prime}})}{\tilde{D}_{f,G}({\bm{X}})}:j^{\prime}\in[L]\Big\}$ $=\Big\{\frac{\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}}+b_{*,j^{\prime}})}{\tilde{D}_{f,\widetilde{G}_{*}}({\bm{X}})}:j^{\prime}\in[L]\Big\},$
@@
$\hskip-100.00015pt=\sum_{j^{\prime}=1}^{L}\exp{(b_{*,j})}\exp((BW_{*,2}W_{*,1j^{\prime}}{\bm{X}})^{\top}{\bm{X}})CW_{*,2}W_{*,1j^{\prime}}{\bm{X}},$ (40)
@@ Appendix C Related Work
Mixture of Experts. Building on classical mixture models with adaptive gating mechanisms (Jacobs et al., 1991; Jordan & Jacobs, 1994; Xu et al., 1994), the MoE model has been extensively studied and refined over the years. Notably, subsequent works (Eigen et al., 2014; Shazeer et al., 2017) introduced the MoE layer as an efficient tool for scaling model capacity. Unlike traditional models that apply uniform parameters to all inputs, the MoE layer applies specific parameter subsets for each input, creating a sparsely activated layer that enables significant capacity growth without proportional increase in computational cost (Fedus et al., 2022; Zhou et al., 2023). This efficiency and scalability have driven its adoption across diverse domains and tasks (Riquelme et al., 2021; Du et al., 2022; Shen et al., 2023).
@@
Theory of Mixture of Experts. Although MoEs have been widely used to scale up large models, their theoretical foundations remain under active development. For example, Ho et al. (2022) focused on input-free gating Gaussian MoEs and showed that under maximum likelihood estimation, the experts’ convergence rates depend on the algebraic independence of the expert functions. Next, Nguyen et al. (2023; 2024a) established convergence rates for both density and parameter estimation in Softmax-gating Gaussian MoEs, linking these rates to the solvability of polynomial systems under Voronoi-based loss functions. More recently, Nguyen et al. (2024b; c) employed least-squares estimation to identify conditions under which expert functions are identifiable. Under these conditions, the resulting estimation rates improve substantially.
@@
Parameter-Efficient Fine-Tuning. Fine-tuning pre-trained foundational models (Dosovitskiy, 2020; Liu et al., 2021; Kirillov et al., 2023) has become a widely adopted strategy for tackling downstream tasks (Xin et al., 2024). While effective, full fine-tuning is resource-intensive, requiring updates to all network parameters and the storage of a separate fine-tuned model for each task (Han et al., 2024). To address these challenges, researchers have increasingly focused on parameter-efficient fine-tuning (PEFT) methods. These methods can be categorized into partial tuning, extra module, and prompt tuning. Partial tuning freezes most of the backbone, fine-tuning only a subset, such as linear heads or a few layers (Mahajan et al., 2018; Chen et al., 2021; He et al., 2022). Extra module methods add trainable parameters to the backbone, such as side structures (Zhang et al., 2020), residual MLP modules in Transformer layers (Houlsby et al., 2019; Cai et al., 2020), or low-rank weight updates (Hu et al., 2021). Despite their promise, partial tuning and extra module methods often face limitations that restrict their applicability. First, they may fail to achieve performance comparable to full fine-tuning (Mahajan et al., 2018; Chen et al., 2021; Jia et al., 2022; He et al., 2023). Second, some methods rely on specific architectural modifications (Rebuffi et al., 2017; Cai et al., 2020; Zhang et al., 2020), limiting their generalizability across different backbone architectures.
@@
Prompting Methods. Prompt tuning (Lester et al., 2021), initially proposed for language tasks, offers a simpler yet effective alternative by introducing learnable parameters to the input sequence of backbone models, updating only these parameters during fine-tuning (Lester et al., 2021; Jia et al., 2022; Le et al., 2025b; a). Despite its simplicity, prompt tuning has demonstrated significant performance gains. Recently, visual prompt tuning has emerged as a promising paradigm in PEFT techniques for computer vision. Current advancements in visual prompt tuning focus on engineering improvements, such as minimizing parameter usage (Han et al., 2023) and broadening applicability to diverse tasks (Yao et al., 2023; Sohn et al., 2023; Yao et al., 2024). However, the theoretical foundations of prompt-based methods remain underexplored. For instance, He et al. (2021) investigated the relationship between prompt tuning and adapter methods, while Le et al. (2024) analyzed these techniques within the context of MoE models. Furthermore, Petrov et al. (2023) highlighted the limitations of prompting, showing that it cannot alter relative attention patterns and instead biases attention layer outputs in a fixed direction.
Adaptive Prompting. A growing body of work investigates adaptive prompting strategies for visual prompt tuning (Zhou et al., 2022). Several approaches employ meta-learning or clustering techniques to tailor prompts to specific downstream datasets. In particular, Huang et al. (2023) and Kim et al. (2023) cluster the downstream data and assign cluster-specific prompts based on the ViT’s initial input representation, ${\bm{X}}^{(0)}$. Similarly, recent prompt-pool methods (Wang et al., 2022; Kim et al., 2024) first forward the input image through an encoder network to obtain a [CLS] representation, which is then used to select or generate prompts at every attention layer. In all these cases, the prompt experts effectively become functions of the first-layer input, ${\bm{X}}^{(0)}$, rather than of the current layer input, ${\bm{X}}^{(l)}$.
Our proposed method, VAPT, instead generates prompts at each layer conditioned on the layer-specific representation, ${\bm{X}}^{(l)}$. This layer-wise adaptation follows directly from our theoretical framework, which establishes a connection between MoE and VPT and implies that both the experts and their gating functions should depend on ${\bm{X}}^{(l)}$. This perspective enables a more principled and analytically grounded examination of model behavior. Importantly, the motivation for VAPT stems from the functional role of prompt experts and their distinction from pre-trained experts, rather than from a purely empirical desire to assign different prompts to different inputs.
Moreover, while some adaptive prompting approaches rely on heuristics or complex image-to-prompt transformations that make their effects difficult to interpret (Huang et al., 2024; Xiao et al., 2025), VAPT adopts a simple yet effective formulation. This design choice leads to the clean expert definitions in Equations (13) and (14), and supports a rigorous theoretical analysis without sacrificing empirical performance—an aspect largely absent in prior work.
Most recently, Zeng et al. (2025) explicitly introduce MoE components into the Transformer architecture, including new routing modules for sparse prompt activation. In contrast, our analysis in Section 3 shows that an implicit MoE structure is already embedded in existing visual prompt tuning frameworks. Consequently, VAPT does not require additional architectural components or routing mechanisms. Instead, it refines the formulation of prompt experts within this inherent implicit MoE structure. In essence, VAPT enhances the expressive capacity of the underlying prompt experts while leaving the backbone architecture unchanged.
@@ Appendix D Implementation Details
Table 4: Specifications of the datasets used in our experiments, which include 24 datasets from two benchmarks: FGVC and VTAB-1K (Zhai et al., 2019).
@@
CUB-200-2011 (Wah et al., 2011) Fine-grained recognition of bird species 200 5,394 600 5,794
NABirds (Van Horn et al., 2015) Fine-grained recognition of bird species 55 21,536 2,393 24,633
Oxford Flowers (Nilsback & Zisserman, 2008) Fine-grained recognition of flower species 102 1,020 1,020 6,149
Stanford Dogs (Khosla et al., 2011) Fine-grained recognition of dog species 120 10,800 1,200 8,580
Stanford Cars (Gebru et al., 2017) Fine-grained recognition of car 196 7,329 815 8,041
@@
CIFAR-100 (Krizhevsky et al., 2009) Natural 100 800/1000 200 10,000
Caltech101 (Fei-Fei et al., 2006) 102 6,084
DTD (Cimpoi et al., 2014) 47 1,880
Flowers102 (Nilsback & Zisserman, 2008) 102 6,149
Pets (Parkhi et al., 2012) 37 3,669
SVHN (Netzer et al., 2011) 10 26,032
Sun397 (Xiao et al., 2010) 397 21,750
Patch Camelyon (Veeling et al., 2018) Specialized 2 800/1000 200 32,768
EuroSAT (Helber et al., 2019) 10 5,400
Resisc45 (Cheng et al., 2017) 45 6,300
Retinopathy (Graham, 2015) 5 42,670
Clevr/count (Johnson et al., 2017) Structured 8 800/1000 200 15,000
Clevr/distance (Johnson et al., 2017) 6 15,000
DMLab (Beattie et al., 2016) 6 22,735
KITTI/distance (Geiger et al., 2013) 4 711
dSprites/loc (Matthey et al., 2017) 16 73,728
dSprites/ori (Matthey et al., 2017) 16 73,728
SmallNORB/azi (LeCun et al., 2004) 18 12,150
SmallNORB/ele (LeCun et al., 2004) 9 12,150
Dataset Specifications. Table 4 provides a summary of the statistics and details of the classification datasets evaluated in this paper. Figure 3 presents image examples from all 24 datasets. Following Jia et al. (2022), each FGVC dataset is randomly split into 90% train and 10% val, with the val set used for hyperparameter tuning. For VTAB-1K, we use an 800-200 split for tuning, and we train on all available data for the final evaluation.
Pre-trained Backbone Specifications. This study explores two primary groups of Vision Transformer (ViT) models. The first group comprises conventional ViT architectures (Dosovitskiy, 2020), including ViT-Base, ViT-Large, and ViT-Huge, pre-trained on the large-scale ImageNet-21K (Deng et al., 2009) dataset. The second group focuses on self-supervised models, specifically Masked Autoencoders (MAE) (He et al., 2022) and Momentum Contrast v3 (MoCo v3) (Chen et al., 2021), both pre-trained on ImageNet-1K. Unless otherwise specified, we use ViT-B/16 with supervised pre-training on ImageNet-21K dataset by default.
@@
Hyperparameters. We use the val set of each dataset to find the optimal prompt length $N_{p}$, kernel size $K$, and hidden dimension $r$ of the feature projector. Following Jia et al. (2022), the search range for $N_{p}$ is $\{1,5,10,50,100,200\}$. Given the relatively small resolution of our feature map ${\bm{X}}_{\mathrm{img}}$ (e.g., $14\times 14$ for ViT), the kernel size $K$ is selected from $\{2,3,4\}$. For the hidden dimension $r$, we explore the range $\{4,8,16,32,64,128,256\}$. To identify the optimal learning rate and weight decay, we perform a grid search, consistent with Mahajan et al. (2018); Jia et al. (2022). The learning rate is searched within $\{50,25,10,5,2.5,1,0.5,0.25,0.1,0.05\}$, while the weight decay is chosen from $\{0.01,0.001,0.0001,0.0\}$. We adopt the batch size settings of Jia et al. (2022), using values of 64 and 128. The models are optimized with SGD for 100 epochs, employing a cosine decay learning rate schedule with 10 warm-up epochs.
Reproducibility. Our method, VAPT, is implemented in PyTorch (Paszke et al., 2019). All experimental workflows, including training and evaluation, were conducted on NVIDIA A100-40GB GPUs. The entire implementation will be made openly available to ensure reproducibility and facilitate future research.